Assume the economy has an MPC of 0.8. An increase in taxes of $20 billion (∆T= 20) would lower the equilibrium level of GDP (- ∆Y=?) by
Answer #1

We know that in equilibrium, Y= c + I+G - [closed Ewnomy] where, Y GDP c→ Consumption expenditure I Investment a > Gouernmenty = thy - 6T + I +G =Y-by = E-bT + I ta - YC6-b) = ē - bT & Ith Now Differentiate with respect to T Assuming everything elsdy = -008 X 20 X109 1-008 16=0.8 d] =$20 billion = dy = - 0,8 x 204109 ► dy --4x20x109 = dy =-80x109 = dy = -80 billion. So,